# Title keyword 5 for citation Title for citation Title for citation

### v1

@@ 4.2.2 Digital Twins for 6G Network native AI
DT technology is pivotal in developing 6G systems and various communication domains. By enabling precise modeling and real-time interaction with physical entities, DT provides practical solutions for optimizing resource allocation, enhancing communication security, and improving service management. For instance, DT optimizes computation offloading in industrial IoT in ultra-reliable, low-latency communication links. Integrating DT with blockchain and federated learning technologies in wireless communication strengthens system security and privacy protection. In the vehicular network domain, DT leverages traffic prediction, edge computing, and load-balancing techniques to enhance real-time data processing and overall system efficiency. DT is driving innovation and development across multiple dimensions of communication systems, demonstrating its transformative potential. The application of DT technology in communication systems is currently experiencing rapid development, showcasing significant potential across various domains, particularly in the research and development of 5G and the emerging 6G networks. In industrial IoT environments, DT technology plays a crucial role. For instance, the authors of [216] address minimizing computation offloading delays in ultra-reliable low-latency communication (URLLC) links by leveraging DT-enabled wireless edge networks. The study focuses on modeling the computational capabilities of edge servers and optimizing system-wide resource allocation.
To enable wireless communication systems based on proactive online learning, the authors of [217] introduce a DT design framework. This framework considers aspects such as twin object access, security and privacy, and air interface design, while also discussing the challenges and opportunities of DT in wireless systems. Additionally, the authors of [218] propose a blockchain-empowered federated learning (FL) framework to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven reinforcement learning to optimize user scheduling and spectrum resource allocation.
@@ 4.2.4 Secure Multi-Party Computation
With the development of 6G networks and AI technology, SMPC provides solid guarantees for privacy protection in 6G networks based on the technologies mentioned above. It has been widely applied in privacy-preserving machine learning and other related areas. Recent research has leveraged SMPC for secure federated learning (FL) model aggregation. FL involves training machine learning models on numerous decentralized devices while localizing the training data. This collaborative approach allows multiple parties to train models without exposing their training data. For example, to counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [241]. To address the issues of high communication overhead and poor scalability in traditional SMPC, a two-phase SMPC-based FL framework has been proposed [242]. This framework enables multiple clients to train AI models while collaboratively protecting their data privacy. To mitigate the high communication costs and poor scalability associated with aggregating global models at only some nodes, a hierarchical FL architecture using SMPC has been developed. This architecture improves communication efficiency and scalability without compromising privacy [243].
@@ 6.1.3 Precision Medical
Healthcare is closely related to individual life and health safety, making the accuracy and reliability of predictive diagnostics crucial. Traditional predictive diagnostics based on single AI models face limitations in accuracy, reliability, and real-time performance due to the maturity of AI algorithms and the availability of case data samples. By leveraging AI over 6G networks, geographically dispersed medical institutions and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances federated learning and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions, thereby increasing trust in the prospects of smart healthcare.
@@ 7 Standardization Process
3GPP has conducted relevant research on network intelligence and automation and introduced network data analytics function (NWDAF) as a core network AI element since Release-15. NWDAF can collect data from any network functions in 5GC and third-party application functions. Due to time limitations, only slice-specific network data analytics is supported in Rel-15. Various network data and third-party application data from AF and 5GC network functions could be used as input to NWDAF, and output analytic results from NWDAF could be used by other NFs to guide network optimization, such as the following use cases been introduced in the Release-16: customized mobile management, 5G quality of service (QoS) enhancement, user plane function (UPF) selection, network functions load balancing, slicing SLA guarantee, etc. Rel-17 allows multiple NWDAFs to be deployed in a distributed architecture. Another enhancement is decomposing the monolithic tWDAF into a Model Training Logic Function, Analytics Logical Function, and Analytics Data Repository Function. Further, the data collection coordination function (DCCF) enhances the data collection framework. Rel-18, the model performance monitoring, multi-vendor ML Model sharing, Horizontal Federated Learning, and six more use cases have been standardized.
@@ 8.2 Challenges of Network For AI
These challenges necessitate a concerted effort from academia and industry to develop AI models with enhanced robustness and generalization capabilities. This involves advancing research in adaptive learning techniques, transfer learning, and federated learning for the academic community to handle better the sporadic and diverse data inherent in 6G networks. For the industry, there is a pressing need to design and implement resilient AI-driven network management systems capable of real-time adaptation to fluctuating conditions. Addressing these challenges satisfactorily will be crucial for leveraging the full potential of AI in optimizing the performance and efficiency of 6G mobile networks.

### v2

@@ 4.2.2 Digital Twins for 6G Network native AI
DT technology is pivotal in developing 6G systems and various communication domains. By enabling precise modeling and real-time interaction with physical entities, DT provides practical solutions for optimizing resource allocation, enhancing communication security, and improving service management. For instance, DT optimizes computation offloading in industrial IoT in ultra-reliable, low-latency communication links. Integrating DT with blockchain and FL technologies in wireless communication strengthens system security and privacy protection. In the vehicular network domain, DT leverages traffic prediction, edge computing, and load-balancing techniques to enhance real-time data processing and overall system efficiency. DT is driving innovation and development across multiple dimensions of communication systems, demonstrating its transformative potential. The application of DT technology in communication systems is currently experiencing rapid development, showcasing significant potential across various domains, particularly in the research and development of 5G and the emerging 6G networks. In industrial IoT environments, DT technology plays a crucial role. For instance, the authors of [215] address minimizing computation offloading delays in ultra-reliable low-latency communication (URLLC) links by leveraging DT-enabled wireless edge networks. The study focuses on modeling the computational capabilities of edge servers and optimizing system-wide resource allocation.
To enable wireless communication systems based on proactive online learning, the authors of [216] introduce a DT design framework. This framework considers aspects such as twin object access, security and privacy, and air interface design, while also discussing the challenges and opportunities of DT in wireless systems. Additionally, the authors of [217] propose a blockchain-empowered FL framework to enhance communication security and data privacy within DT-enabled edge networks. The study incorporates asynchronous aggregation mechanisms and DT-driven reinforcement learning to optimize user scheduling and spectrum resource allocation.
@@ 4.2.4 Secure Multi-Party Computation
With the development of 6G networks and AI technology, SMPC provides solid guarantees for privacy protection in 6G networks based on the technologies mentioned above. It has been widely applied in privacy-preserving machine learning and other related areas. Recent research has leveraged SMPC for secure FL model aggregation. FL involves training machine learning models on numerous decentralized devices while localizing the training data. This collaborative approach allows multiple parties to train models without exposing their training data. For example, to counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [240]. To address the issues of high communication overhead and poor scalability in traditional SMPC, a two-phase SMPC-based FL framework has been proposed [241]. This framework enables multiple clients to train AI models while collaboratively protecting their data privacy. To mitigate the high communication costs and poor scalability associated with aggregating global models at only some nodes, a hierarchical FL architecture using SMPC has been developed. This architecture improves communication efficiency and scalability without compromising privacy [242].
@@ 6.1.3 Precision Medical
Healthcare is closely related to individual life and health safety, making the accuracy and reliability of predictive diagnostics crucial. Traditional predictive diagnostics based on single AI models face limitations in accuracy, reliability, and real-time performance due to the maturity of AI algorithms and the availability of case data samples. By leveraging AI over 6G networks, geographically dispersed medical institutions and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances FL and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions, thereby increasing trust in the prospects of smart healthcare.
@@ 7 Standardization Process
3GPP has conducted relevant research on network intelligence and automation and introduced network data analytics function (NWDAF) as a core network AI element since Release-15. NWDAF can collect data from any network functions in 5GC and third-party application functions. Due to time limitations, only slice-specific network data analytics is supported in Rel-15. Various network data and third-party application data from AF and 5GC network functions could be used as input to NWDAF, and output analytic results from NWDAF could be used by other NFs to guide network optimization, such as the following use cases been introduced in the Release-16: customized mobile management, 5G quality of service (QoS) enhancement, user plane function (UPF) selection, network functions load balancing, slicing SLA guarantee, etc. Rel-17 allows multiple NWDAFs to be deployed in a distributed architecture. Another enhancement is decomposing the monolithic tWDAF into a Model Training Logic Function, Analytics Logical Function, and Analytics Data Repository Function. Further, the data collection coordination function (DCCF) enhances the data collection framework. Rel-18, the model performance monitoring, multi-vendor ML Model sharing, Horizontal FL, and six more use cases have been standardized.
@@ 8.2 Challenges of Network For AI
These challenges necessitate a concerted effort from academia and industry to develop AI models with enhanced robustness and generalization capabilities. This involves advancing research in adaptive learning techniques, transfer learning, and FL for the academic community to handle better the sporadic and diverse data inherent in 6G networks. For the industry, there is a pressing need to design and implement resilient AI-driven network management systems capable of real-time adaptation to fluctuating conditions. Addressing these challenges satisfactorily will be crucial for leveraging the full potential of AI in optimizing the performance and efficiency of 6G mobile networks.